OpenStudy (anonymous):

EASY QUESTION + MEDAL AND FAVORITE: Given that f(x) = 3x + 1 and g(x) = the quantity of 4x plus 2 divided by 3, solve for g(f(0)).

OpenStudy (anonymous):

f(x) = 3x + 1 g(x) = \[4x + 2 \over 3\] first find f(0) f(x) = 3x + 1 F(0) = 3(0) + 1 0 = 0 + 1 0 = 0? f(0) = 1 plug that in g(f(0)) g(1) g(1).. find it 4x + 2 / 3 4(1) + 2 / 3 4 + 2 = 6 6/3 2